OpenStudy (anonymous):

Math Analysis: cos theta*cot(90-theta)=?

OpenStudy (anonymous):

sin(theta)

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

cot(90-x)=tan(x)=sin(x)/cos(x) =>cos(x)*cot(90-x)=cos(x)*sin(x)/cos(x)=sin(x)

OpenStudy (anonymous):

okay thx xD